Two electrons are 10^-5 m apart. One electron is fixed on its position. The other, originally at rest, is released. How fast is the free electron moving when it is 10^−3 m from the fixed charge? Apply conservation of energy.


1
Expert's answer
2022-04-15T16:32:04-0400

"kq^2(\frac 1{r_1}-\frac 1{r_2})=\frac{mv^2}2,"

"v=\sqrt{\frac{2kq^2(r_2-r_1)}{mr_1r_2}}=2.25\cdot 10^7~\frac ms ."